# Beyond Collision Cones: Dynamic Obstacle Avoidance for Nonholonomic Robots via Dynamic Parabolic Control Barrier Functions

### v1

@@ I INTRODUCTION
This paper introduces a Dynamic Parabolic Control Barrier Function (DPCBF) that explicitly incorporates both clearance and the magnitude of the relative velocity. Instead of a fixed cone, we define a state-dependent parabolic safety boundary whose curvature and vertex adapt with distance and relative velocity (see Fig. 1b). This design allows a less conservative safety constraints, improving the CBF-based controller’s feasibility in cluttered, dynamic environments. The main contributions of this work are:
@@ II-A Control Barrier Functions
([11, Theorem 1]) Let $h$ satisfy the CBF condition (3) and define
@@ II-B Bicycle Model
In this paper, we consider a robot modeled by the kinematic bicycle model [12, 13] (see Fig. 2). The state is ${\bm{x}}=[x,y,\theta,v]^{\top}$, where $x,y$ are the position of the vehicle’s center of mass (CoM), $\theta$ is the heading angle, and $v$ is the forward velocity. The control inputs are longitudinal acceleration $a$ and forward wheel steering angle $\delta$. Let $\ell_{f}$ and $\ell_{r}$ denote the distances from the CoM to the front and rear axles, respectively, and define the slip angle $\beta=\tan^{-1}(\tan(\delta)\,\ell_{r}/(\ell_{f}+\ell_{r}))$.
Figure 2: Schematic of the kinematic bicycle model. The robot’s state defined by its Center of Mass (CoM) position $(x,y)$, heading angle $\theta$, and forward velocity $v$. The distances from the CoM to the front and rear axles are $\ell_{f}$ and $\ell_{r}$, respectively. The front wheel steering angle is $\delta$, and $\beta$ is the resulting vehicle slip angle.
@@ II-C Obstacle Model
We model a scenario with multiple dynamic obstacles. The state of the $j$-th dynamic obstacle, $j=\{1,\ldots,N_{\textup{obs}}\}$, is represented by
@@
The dynamics of the $j$-th obstacle is described by a unicycle model with constant velocity: $\dot{x}_{\textup{obs}}^{j}=v_{\textup{obs}}^{j}\cos\theta_{\textup{obs}}^{j}$, $\dot{y}_{\textup{obs}}^{j}=v_{\textup{obs}}^{j}\sin\theta_{\textup{obs}}^{j}$. We assume the state of the obstacles are fully observable. For the remainder of the paper, we omit the superscript $j$ and describe the CBF constraint for each obstacle for notational simplicity.
II-D Distanced-Based CBF
@@
Because it only considers distance, this barrier is not a CBF in a general, except for simple systems where control inputs directly affect velocity. Its myopic nature makes it particularly unsuitable for systems with nonholonomic constraints [9].
@@ II-E Collision Cone CBF
Figure 3: A closer look on collision cone-based CBF. (a) Single obstacle: if $h_{\textup{cc}}({\bm{x}}(t_{0}))\geq 0$, the CBF constraint keeps $h_{\textup{cc}}({\bm{x}}(t))\geq 0$ for all $t\geq t_{0}$. (b) Five obstacles: if $h_{j,\textup{cc}}({\bm{x}}(t))\geq 0$ for all $j$-th obstacle and the CBF-QP is feasible at $t$, safety is at least maintained at a given time $t$. (c) Even with $h_{j,\textup{cc}}({\bm{x}}(t_{0}))\geq 0$ for all $j$-th obstacle, cone intersections can leave no admissible relative velocity direction, leading the CBF-QP infeasible. (d) Given the initial configuration where the robot is surrounded by the union of the collision cones, there is no feasible solution to the CBF-QP even though a large collision-free area exists nearby.
@@ III DYNAMIC PARABOLIC CBF
In this paper, we present a novel CBF formulation for dynamic obstacle collision avoidance tasks. The existing methods are either not safe with non-static obstacles, or overly prone to be infeasible when multiple obstacles are nearby. Therefore, we focus on improving the key criteria: (i) guarantee safety for dynamic obstacles under input constraints, and (ii) improve feasibility of the resulting CBF-based controller.
@@ III-A1 Relative Coordinates
We refer to the rotated frame by the rotation matrix $\mathbf{R}$ as Line-of-Sight (LoS) frame throughout this paper (see Fig. 4b). Finally, we define the relative velocity in the LoS frame:
@@ III-A2 CBF Formulation and Design Maps
Let $r\in\mathbb{R}$ be the combined radius of robot and obstacle define in (8). Then we have
@@
We propose Dynamic Parabolic CBF (DPCBF) as the following:
@@
where $\tilde{v}_{\textup{rel},x}$ and $\tilde{v}_{\textup{rel},y}$ are the relative velocities between the robot and the obstacle in the LoS frame. Since $\lambda({\bm{x}})$ and $\mu({\bm{x}})$ are functions of the relative distance and speed, the parabola representing the unsafe set dynamically adapts its shape online to the current situation. The CBF is now defined by measuring how close the endpoint of the rotated relative velocity (12) is from a specific parabolic region in this new plane (see Fig. 5):
@@
This provides a significant advantage over cone-based methods. As illustrated in Fig. 1b, the boundary of the unsafe set (16) does not intersect the robot’s current position $(x,y)$ whenever the clearance to the obstacle, $d({\bm{x}})$, is non-zero. This shift creates a feasible space where motion toward an obstacle is no longer treated immediately as unsafe. Instead, safety is now evaluated jointly on the current clearance $d({\bm{x}})$ and the relative velocity ${\bm{v}}_{\textup{rel}}$. We empirically show in Sec. IV that this design promotes improved feasibility of the DPCBF-based controller compared to prior works.
Figure 5: Three examples describe how (16) shapes the safety boundary. (a) When $k_{\lambda}=0$ ($k_{\mu}$ being active), as the obstacle gets closer, the parabola’s vertex moves toward the robot, shrinking the safe region. (b) Where $k_{\mu}=0$ ($k_{\lambda}$ being active), as the obstacle approaches or the relative velocity magnitude increases, the curvature of the parabola decrease, leading a larger unsafe set. (c) In DPCBF, where $k_{\lambda}\neq 0$ and $k_{\mu}\neq 0$, both the vertex and the curvature of the parabola adapts dynamically.
@@ III-B Validity of DPCBF
$\big|C^{a}({\bm{x}})\big|=\Bigg|\underbrace{\Bigl[-1+k_{\lambda}\frac{d({\bm{x}})}{\|v_{\textup{rel}}\|^{3}}\,v_{\textup{obs}}\,\cos\tilde{\theta}_{\textup{obs}}\tilde{v}_{\textup{rel},y}^{2}\Bigr]}_{\coloneqq\eta_{a,\cos}({\bm{x}})}\,\cos\tilde{\theta}$
$+\underbrace{\Bigl[k_{\lambda}\frac{d({\bm{x}})}{\|v_{\textup{rel}}\|^{3}}v_{\textup{obs}}\sin\tilde{\theta}_{\textup{obs}}\tilde{v}_{\textup{rel},y}^{2}-2k_{\lambda}\frac{d({\bm{x}})}{\|v_{\textup{rel}}\|}\tilde{v}_{\textup{rel},y}\Bigr]}_{\coloneqq\eta_{a,\sin}({\bm{x}})}\sin\tilde{\theta}$
$+\underbrace{\Bigl[-k_{\lambda}\frac{d({\bm{x}})}{\|v_{\textup{rel}}\|^{3}}v\,\tilde{v}_{\textup{rel},y}^{2}\Bigr]}_{\coloneqq\eta_{a,0}({\bm{x}})}\Bigg|,$ (18)
@@
$\inf_{{\bm{x}}\in\partial\mathcal{C}_{1}}\frac{L_{f}h({\bm{x}})}{v}+\inf_{{\bm{x}}\in\partial\mathcal{C}_{1}}\biggr|\frac{C^{a}({\bm{x}})}{v}\biggl|a_{\max}$
$\qquad\qquad\qquad\qquad\quad+\inf_{{\bm{x}}\in\partial\mathcal{C}_{1}}\biggr|\frac{C^{\beta}({\bm{x}})}{v}\biggl|\beta_{\max}\geq 0$ (25)
@@
Figure 6: Demonstration of the proposed DPCBF’s navigation behavior in a surrounded environment with ten dynamic obstacles. Note that for the same configuration, collision cone-based methods are infeasible, as was demonstrated in Fig 3d. (Left) At $t=0$ s, with an $v(t_{0})=0.5$ m/s, despite being surrounded by the unsafe sets, the QP with DPCBF constraints finds a feasible solution by ensuring the robot’s relative velocity vectors lie outside the dynamic parabolic boundaries, allowing it to proceed safely. (Center) By $t=4.4$ s, the robot successfully maneuvers through a narrow passage. This is possible due to the less conservative formulation of DPCBF, which provides the necessary control flexibility in confined spaces. (Right) The robot safely navigates through the obstacles and reaches the goal at $t=7.9$ s.
@@
As is common in CBF analysis, including the prior works we evaluate against, the safety guarantee in Theorem 1 holds for a single CBF constraint, corresponding to one obstacle. For methods on composing multiple CBF constraints into a single constraint, we refer the readers to [21, 6]. A formal investigation into the composition of multiple DPCBFs under input constraints is outside the scope of this paper. We evaluate the performance of our DPCBF-based controller against compared methods using a QP with multiple constraints in Sec. IV.
@@ IV-C Experimental Results
Performance Analysis in Dense Dynamic Environments. To test the core hypothesis that DPCBF alleviate infeasibility issues while ensuring safety, we simulate navigation in environments with an increasing number of dynamic obstacles, from 1 to 100. The results are summarized in Fig. 8. We first evaluate the methods in single-obstacle scenarios, where the formal safety guarantee holds for all methods except for Dynamic zone-based CBF. As expected from the theoretical guarantee, they achieved a 100% success rate. Dynamic zone-based CBF resulted in a 1.7% infeasibility rate given that it is not a valid CBF for the kinematic bicycle model, regardless of the number of constraints. The performance of the compared methods drops dramatically as the number of dynamic obstacles increases, resulting in frequent QP infeasibility or even collisions. Notably, DPCBF achieves a 100% success rate even in the 10-obstacle cases. This shows that the collision-cone based methods [15, 22] suffer in obstacle-dense environments, where overlapping collision cones severely constrain the feasible control space, leading to frequent QP failures.
@@
Figure 10: QP cost and resulting control inputs over time for the dense scenario from Fig. 7. The peaks in QP cost near $t=21$ s correspond to the most obstacle-dense moments of the trajectory as shown in Fig. 7.
Qualitative Trajectory Analysis. Fig. 7 visualizes a challenging navigation scenario where the DPCBF-based QP controlling the robot to navigate through dense group of 100 dynamic obstacles with a maximum obstacle radius of $r_{\textup{obs},\max}=0.7$ m. We also visualize the velocity of the kinematic bicycle along its trajectory, with the corresponding QP cost and control inputs shown in Fig. 10. This demonstrates that DPCBF constraints guide the CBF-QP to effectively adjust both longitudinal and lateral motion around multiple obstacles, successfully performing safe navigation. Importantly, at snapshots taken at $t=10$ s and $t=21$ s, the union of the unsafe sets does not render the feasible space empty, whereas methods based on VO or collision cone would lead to infeasibility in the same configurations. The robot is also able to regain high velocity at $t=33$ s when the obstacles are no longer driving towards it. These examples highlight how DPCBF actively modifies the nominal control inputs to guarantee safety without being overly conservative.

### v2

@@ I INTRODUCTION
This paper introduces a Dynamic Parabolic Control Barrier Function (DPCBF) that explicitly incorporates both clearance and the magnitude of the relative velocity. Instead of a fixed cone, we define a state-dependent parabolic safety boundary whose curvature and vertex adapt with distance and relative velocity (see Fig. 1b). This design yields less conservative safety constraints, improving the feasibility of the CBF-based controller in cluttered, dynamic environments. The main contributions of this work are:
@@ II-A Control Barrier Functions
([1, Theorem 1]) Let $h$ satisfy the CBF condition (3) and define
@@ II-B Bicycle Model
In this paper, we consider a robot modeled by the kinematic bicycle model [13, 9] (see Fig. 2). The state is ${\bm{x}}=[x,y,\theta,v]^{\top}$, where $(x,y)$ denotes the position of the vehicle’s center of mass (CoM), $\theta$ is the heading angle, and $v$ is the forward velocity. The control inputs are longitudinal acceleration $a$ and the forward-wheel steering angle $\delta$. Let $\ell_{f}$ and $\ell_{r}$ denote the distances from the CoM to the front and rear axles, respectively, and define the slip angle $\beta=\tan^{-1}\bigl(\tan(\delta)\,\ell_{r}/(\ell_{f}+\ell_{r})\bigr)$.
Figure 2: Schematic of the kinematic bicycle model. The robot’s state is defined by its Center of Mass (CoM) position $(x,y)$, heading angle $\theta$, and forward velocity $v$. The distances from the CoM to the front and rear axles are $\ell_{f}$ and $\ell_{r}$, respectively. The front-wheel steering angle is $\delta$, and $\beta$ is the resulting vehicle slip angle.
@@ II-C Obstacle Model
We model a scenario with multiple dynamic obstacles. The state of the $j$-th dynamic obstacle, where $j\in\{1,\ldots,N_{\textup{obs}}\}$, is represented by
@@
The dynamics of the $j$-th obstacle is described by a unicycle model with constant velocity: $\dot{x}_{\textup{obs}}^{j}=v_{\textup{obs}}^{j}\cos\theta_{\textup{obs}}^{j}$ and $\dot{y}_{\textup{obs}}^{j}=v_{\textup{obs}}^{j}\sin\theta_{\textup{obs}}^{j}$. We assume that the obstacle states are fully observable. For the remainder of the paper, we omit the superscript $j$ and describe the CBF constraint for each obstacle for notational simplicity.
II-D Distance-Based CBF
@@
Because it only considers distance, this barrier is not, in general, a CBF except for simple systems in which the control inputs directly affect the velocity. Its myopic nature makes it particularly unsuitable for systems with nonholonomic constraints [17].
@@ II-E Collision Cone CBF
Figure 3: A closer look at collision cone-based CBF. (a) Single obstacle: if $h_{\textup{cc}}({\bm{x}}(t_{0}))\geq 0$, the CBF constraint keeps $h_{\textup{cc}}({\bm{x}}(t))\geq 0$ for all $t\geq t_{0}$. (b) Five obstacles: if $h_{j,\textup{cc}}({\bm{x}}(t))\geq 0$ for all $j$-th obstacle and the CBF-QP is feasible at $t$, safety is at least maintained at a given time $t$. (c) Even with $h_{j,\textup{cc}}({\bm{x}}(t_{0}))\geq 0$ for all $j$-th obstacle, cone intersections can leave no admissible relative velocity direction, leading the CBF-QP infeasible. (d) Given the initial configuration where the robot is surrounded by the union of the collision cones, there is no feasible solution to the CBF-QP even though a large collision-free area exists nearby.
@@ III DYNAMIC PARABOLIC CBF
In this paper, we present a novel CBF formulation for dynamic-obstacle collision avoidance tasks. Existing methods either do not provide safety guarantees for moving obstacles or are prone to infeasibility when multiple obstacles are nearby. Accordingly, we focus on improving the key criteria: (i) guaranteeing safety for dynamic obstacles under input constraints, and (ii) improving the feasibility of the resulting CBF-based controller.
@@ III-A1 Relative Coordinates
We refer to the rotated frame defined by the rotation matrix $\mathbf{R}$ as Line-of-Sight (LoS) frame throughout this paper (see Fig. 4b). Finally, we define the relative velocity in the LoS frame:
@@ III-A2 CBF Formulation and Design Maps
Let $r\in\mathbb{R}$ be the combined robot-obstacle radius defined in (8). Then,
@@
We propose the Dynamic Parabolic CBF (DPCBF) as follows:
@@
where $\tilde{v}_{\textup{rel},x}$ and $\tilde{v}_{\textup{rel},y}$ are the components of the relative velocity between the robot and the obstacle in the LoS frame. Since $\lambda({\bm{x}})$ and $\mu({\bm{x}})$ depend on the relative distance and speed, the parabola representing the unsafe set dynamically adapts its shape online to the current situation. The CBF is now defined by measuring how close the endpoint of the rotated relative velocity (12) is to a specific parabolic region in this new plane (see Fig. 5):
@@
This provides a significant advantage over cone-based methods. As illustrated in Fig. 1b, the boundary of the unsafe set (16) does not pass through the origin of the LoS-frame relative-velocity plane whenever the clearance to the obstacle, $d({\bm{x}})$, is non-zero. This shift creates a feasible space where motion toward an obstacle is no longer treated immediately as unsafe. Instead, safety is now evaluated jointly using the current clearance $d({\bm{x}})$ and the relative velocity ${\bm{v}}_{\textup{rel}}$. We empirically show in Sec. IV that this design improves the feasibility of the DPCBF-based controller compared with prior work.
Figure 5: Three examples illustrating how (16) shapes the safety boundary. (a) When $k_{\lambda}=0$ (with $k_{\mu}$ active), as the obstacle gets closer, the parabola’s vertex moves toward the robot, shrinking the safe region. (b) When $k_{\mu}=0$ (with $k_{\lambda}$ active), as the obstacle approaches or the relative velocity magnitude increases, the curvature of the parabola decrease, leading to a larger unsafe set. (c) In DPCBF, where $k_{\lambda}\neq 0$ and $k_{\mu}\neq 0$, both the vertex and the curvature of the parabola adapt dynamically.
@@ III-B Validity of DPCBF
$\big|C^{a}({\bm{x}})\big|=\Bigg|\underbrace{\Bigl[-1+k_{\lambda}\frac{d({\bm{x}})}{\|{\bm{v}}_{\textup{rel}}\|^{3}}\,v_{\textup{obs}}\,\cos\tilde{\theta}_{\textup{obs}}\tilde{v}_{\textup{rel},y}^{2}\Bigr]}_{\coloneqq\eta_{a,\cos}({\bm{x}})}\,\cos\tilde{\theta}$
$+\underbrace{\Bigl[k_{\lambda}\frac{d({\bm{x}})}{\|{\bm{v}}_{\textup{rel}}\|^{3}}v_{\textup{obs}}\sin\tilde{\theta}_{\textup{obs}}\tilde{v}_{\textup{rel},y}^{2}-2k_{\lambda}\frac{d({\bm{x}})}{\|{\bm{v}}_{\textup{rel}}\|}\tilde{v}_{\textup{rel},y}\Bigr]}_{\coloneqq\eta_{a,\sin}({\bm{x}})}\sin\tilde{\theta}$
$+\underbrace{\Bigl[-k_{\lambda}\frac{d({\bm{x}})}{\|{\bm{v}}_{\textup{rel}}\|^{3}}v\,\tilde{v}_{\textup{rel},y}^{2}\Bigr]}_{\coloneqq\eta_{a,0}({\bm{x}})}\Bigg|,$ (18)
@@
$\inf_{{\bm{x}}\in\partial\mathcal{C}_{1}}\frac{L_{f}h({\bm{x}})}{v}+\inf_{{\bm{x}}\in\partial\mathcal{C}_{1}}\biggl|\frac{C^{a}({\bm{x}})}{v}\biggr|a_{\max}$
$\qquad\qquad\qquad\qquad\quad+\inf_{{\bm{x}}\in\partial\mathcal{C}_{1}}\biggl|\frac{C^{\beta}({\bm{x}})}{v}\biggr|\beta_{\max}\geq 0$ (25)
@@
Figure 6: Demonstration of the proposed DPCBF’s navigation behavior in a surrounded environment with ten dynamic obstacles. For the same configuration, collision cone-based methods are infeasible, as shown in Fig 3d. (Left) At $t=0$ s, with $v(t_{0})=0.5$ m/s, the QP with DPCBF constraints finds a feasible solution even though the robot is surrounded by unsafe sets, by keeping the relative-velocity vectors outside the dynamic parabolic boundaries, allowing it to proceed safely. (Center) By $t=4.4$ s, the robot successfully maneuvers through a narrow passage. This is possible due to the less conservative formulation of DPCBF, which provides the necessary control flexibility in confined spaces. (Right) The robot safely navigates through the obstacles and reaches the goal at $t=7.9$ s.
@@
As is common in CBF analysis, including the prior works we evaluate against, the safety guarantee in Theorem 1 holds for a single CBF constraint, corresponding to one obstacle. For methods on composing multiple CBF constraints into a single constraint, we refer the readers to [3, 6]. A formal investigation into the composition of multiple DPCBFs under input constraints is outside the scope of this paper. We evaluate the performance of our DPCBF-based controller against the compared methods using a QP with multiple constraints in Sec. IV.
@@ IV-C Experimental Results
Performance Analysis in Dense Dynamic Environments. To test the core hypothesis that DPCBF alleviates infeasibility issues while ensuring safety, we simulate navigation in environments with an increasing number of dynamic obstacles, from 1 to 100. The results are summarized in Fig. 8. We first evaluate the methods in single-obstacle scenarios, where the formal safety guarantee holds for all methods except for Dynamic zone-based CBF. As expected from the theoretical guarantee, all such methods achieve a 100% success rate. The Dynamic zone-based CBF exhibits a 1.7% infeasibility rate because it is not a valid CBF for the kinematic bicycle model, regardless of the number of constraints. The performance of the compared methods drops dramatically as the number of dynamic obstacles increases, resulting in frequent QP infeasibility or even collisions. Notably, DPCBF achieves a 100% success rate even in the 10-obstacle cases. This shows that the collision-cone based methods [18, 14] suffer in obstacle-dense environments, where overlapping collision cones severely constrain the feasible control space, leading to frequent QP failures.
@@
Figure 10: QP cost and resulting control inputs over time for the dense scenario from Fig. 7. The peaks in QP cost near $t=21$ s correspond to the most obstacle-dense moments of the trajectory shown in Fig. 7.
Qualitative Trajectory Analysis. Fig. 7 visualizes a challenging navigation scenario in which the DPCBF-based QP guides the robot through a dense group of 100 dynamic obstacles with a maximum obstacle radius of $r_{\textup{obs},\max}=0.7$ m. We also visualize the velocity of the kinematic bicycle along its trajectory, with the corresponding QP cost and control inputs shown in Fig. 10. This demonstrates that DPCBF constraints guide the CBF-QP to effectively adjust both longitudinal and lateral motion around multiple obstacles, successfully performing safe navigation. Importantly, at snapshots taken at $t=10$ s and $t=21$ s, the union of the unsafe sets does not render the feasible space empty, whereas methods based on VO or collision cone would be infeasible in the same configurations. The robot is also able to regain high velocity at $t=33$ s when the obstacles are no longer driving towards it. These examples highlight how DPCBF actively modifies the nominal control inputs to guarantee safety without being overly conservative.